# Method for simulating open-system dynamics using mid-circuit measurements on a quantum computer

### v1

@@ Abstract
We present a method for simulating the dynamics of an open electronic system on a quantum computer. This approach entails mid-circuit measurements and resets to simulate the addition or removal of fermions from the system and provides a way to apply non-reversible operations to a quantum computer. Using this method, we simulate the dynamics of an open electronic system consisting of a fermionic chain positioned between two conductive leads on the $ibm\_torino$ quantum computer.
@@ I Introduction
Quantum computing shows promise for simulating electronic systems. This is due to the straightforward mapping between the unitary operators that govern electron dynamics and those describing the quantum logic gates that are executed on quantum computers. Consequently, there are many studies that report simulations of closed-system electron dynamics on quantum computers Peruzzo et al. (2014); McClean et al. (2016); O’Malley et al. (2016); Kandala et al. (2019); Wang et al. (2019); Bian et al. (2019); Arute et al. (2020); McArdle et al. (2020); Head-Marsden et al. (2021); Kim et al. (2023). Conversely, it is non-trivial to simulate the dynamics of open quantum systems on a quantum computer. This is largely due to the fact that an open quantum system can have non-reversible dynamics while quantum logic gates are necessarily reversible. Therefore, there are fewer reports of open-system electron dynamics simulations on quantum computers Wei et al. (2018); Sweke et al. (2015); Di Candia et al. (2015); Wei et al. (2016); Wang et al. (2013, 2011); Hu et al. (2020).
@@ II.1 Closed electronic systems
where $\Delta\hat{U}=\prod_{s}e^{-i\hat{P}_{s}t/N_{t}}$, $N_{t}$ is the number of time-steps and $\hat{P}_{s}$ are Pauli strings corresponding to terms in the Hamiltonian. The Lie–Trotter approximation is needed because the $\hat{P}_{s}$ do not commute in general. The operator $\Delta\hat{U}$ can be implemented using quantum logic gates Mansky et al. (2023). This procedure is well-established and has been applied to fermionic simulations on quantum computers Abrams and Lloyd (1997); Whitfield et al. (2011); Stenger et al. (2023) . Simulating the dynamics of open systems, however, is less straightforward and demands a strategy to implement non-reversible operations.
@@ II.2 Open systems
Consider an open fermionic system coupled to metallic contacts. An example of such a system is a chain of spinless fermions between source and drain contacts as illustrated in Fig. 1a. We represent each site by a qubit, where the end qubits are connected to the contacts as shown in Fig. 1b. Transport is modeled by evolution of the system state, in which fermions are injected or removed at the contacts. To simulate the dynamics we introduce the state |$\ket{\Phi(t)}$ at time $t$, starting from the initial state |$\ket{\Phi(0)}=\ket{\Phi_{0}}$.
@@
At each end site and time-point, we probabilistically select between three possible actions: (1) inject a fermion, (2) remove a fermion, (3) do nothing. Consider a quantum register with $L$ qubits and let |$\ket{0}_{q}$ and |$\ket{1}_{q}$ be the basis states for Hilbert space of qubit $q$. The actions are performed through mid-circuit measurements. An electron is injected by setting the qubit measured in |$\ket{0}_{q}$ to |$\ket{1}_{q}$ and removed by setting the qubit measured in |$\ket{1}_{q}$ to |$\ket{0}_{q}$. For each end qubit $q$ and conductor $\alpha$, we define
@@
where $P^{\text{in}}_{q\alpha}$ and $P^{\text{out}}_{q\alpha}$ are the probabilities of electron injection and removal, respectively, $f$ is the Fermi–Dirac distribution function, $\mu_{\alpha}$ is the electrochemical potential of contact $\alpha$, and $\eta_{q\alpha}=\Gamma_{q\alpha}t/N_{t}$, where $\Gamma_{q\alpha}$ is the coupling rate between qubit $q$ and conductor $\alpha$. Note that $N_{t}$ must be sufficiently large such that the probability of fermion injection or removal is less than or equal to one, $N_{t}\geq(\Gamma_{q\alpha}t)^{-1}$ for all $\Gamma_{q\alpha}$. In the limit $N_{t}\rightarrow\infty$, this method simulates the dynamics of a system in contact with two conductors. See Appendix A for a comparison of our method with the Lindblad equation Brasil et al. (2013).
@@ III.1 Model System
We now demonstrate our method for the chain of interacting spinless fermions, described by the Hamiltonian,
$\hat{H}=\gamma\sum_{<i,j>}(\hat{c}^{\dagger}_{i}\hat{c}_{j}+\hat{c}^{\dagger}_{j}\hat{c}_{i})+v\sum_{<i,j>}\hat{n}_{i}\hat{n}_{j},$ (4)
@@
$\begin{split}\hat{H}=\gamma\sum_{<q,p>}(\hat{X}_{q}\hat{X}_{p}+\hat{Y}_{q}\hat{Y}_{p})+v\sum_{<q,p>}(\hat{I}+\hat{Z}_{q})(\hat{I}+\hat{Z}_{p}).\\ \end{split}$ (5)
@@
We then simulate the fermion chain both as a closed system, and as an open system coupled to contacts. In the closed-system case, we set $\eta_{q\alpha}=0$ for all $q$ and $\alpha\in\{\text{S},\text{D}\}$. In the open-system case, we couple the system to a source $f(\mu_{\text{S}})=1$ at the first site $\eta_{1\text{S}}\neq 0$ and a drain $f(\mu_{\text{D}})=0$ at last site $\eta_{L\text{D}}\neq 0$, where $i=L$ is the last site.
@@ III.3 Open-system dynamics
We now compare the dynamics of the open system simulated on a classical computer to those simulated on a quantum computer. Figure 3 shows the dynamics of a seven-site system as simulated on a classical (a-b) and quantum computer (c-d). We consider $\gamma$=3.0 meV and 5.0 meV, and set $v$=10.0 meV, $\eta_{\text{qc}}=0.5$ meV, and $N_{t}/t=2\leavevmode\nobreak\ \text{meV}/\hbar$, in all cases. For all simulations shown in Fig.  3, the density is initially $n_{0}(0)=1.0$ and corresponds to the injection of an electron into the chain. Over time, the densitys at sites $i>0$ increase, as electron density flows towards the drain as indicated by the dotted green line in Fig. 3a. Blue arrows at $i=1$ at correspond to the injection of an electron while those at $i=L$ correspond to the removal of electrons. We see that injecting an electron causes a sharp increase in density while the removal of an electron causes a sharp decrease in density, as expected. When electrons are not immediately removed, the can reflect off of the right end of the wire. Backpropagation of electron density can be seen as marked by the dotted yellow line. This results in a peak in density at the left end not corresponding to the injection of an electron. The flow of density from $i=1$ to $i=7$ is faster for $\gamma$=5.0 meV (Fig. 3b,d) than for $\gamma$=3.0 meV (Fig. 3a,c), on both the classical and quantum computers.
@@ Appendix A Relationship to the Lindblad Equation
where $\hat{H}$ is the Hamiltonian of the system, $\hat{L}_{0}=\sqrt{\Gamma f(\mu)}\hat{c}^{\dagger}_{l}$ is the Lindblad operator for the incoming electrons, $\hat{L}_{1}=\sqrt{\Gamma[1-f(\mu)]}\hat{c}_{l}$ is the Lindblad operator for the outgoing electrons, $\Gamma$ is the coupling rate between the contact and the system, $f(\mu)$ is the Fermi-Dirac distribution function evaluated at the chemical potential of the lead given that the lead has an electrochemical potential $\mu$, $\hat{c}^{\dagger}_{i}$ is the electron creation operator at point $i$, and $l$ is the point of contact. We use selective mid-circuit measurements to reproduce these dynamics on a quantum computer.

### v2

@@ Abstract
We present a method for simulating the dynamics of an open electronic system on a quantum computer. This approach entails mid-circuit measurements and resets to simulate the addition or removal of electrons from the system. Our method provides a way to apply non-reversible operations to a quantum computer without the need for additional qubits. Using this method, we simulate the dynamics of an open electronic system consisting of a chain of electrons positioned between two conductive leads on the $ibm\_torino$ quantum computer. We expect the method to be generally applicable to open systems.
@@ I Introduction
Quantum computing shows promise for simulating electronic systems. This is due to the straightforward mapping between the unitary operators that govern electron dynamics and those describing the quantum logic gates that are executed on quantum computers. Consequently, there are many studies that report simulations of closed-system electron dynamics on quantum computers Peruzzo et al. (2014); McClean et al. (2016); O’Malley et al. (2016); Kandala et al. (2019); Wang et al. (2019); Bian et al. (2019); Arute et al. (2020); McArdle et al. (2020); Head-Marsden et al. (2021); Kim et al. (2023); Wang et al. (2024). Conversely, it is non-trivial to simulate the dynamics of open quantum systems on a quantum computer. This is largely due to the fact that an open quantum system can have non-reversible dynamics while quantum logic gates are necessarily reversible. Therefore, there are fewer reports of open-system electron dynamics simulations on quantum computers Wei et al. (2018); Sweke et al. (2015); Di Candia et al. (2015); Wei et al. (2016); Wang et al. (2013, 2011); Hu et al. (2020); Bian et al. (2023); Li et al. (2023).
@@ II.1 Closed electronic systems
where $\Delta\hat{U}=\prod_{s}e^{-i\hat{P}_{s}t/N_{t}}$, $N_{t}$ is the number of time-steps and $\hat{P}_{s}$ are Pauli strings corresponding to terms in the Hamiltonian. The Lie–Trotter approximation is needed because the $\hat{P}_{s}$ do not commute in general. The operator $\Delta\hat{U}$ can be implemented using quantum logic gates Mansky et al. (2023). This procedure is well-established and has been applied to electronic simulations on quantum computers Abrams and Lloyd (1997); Whitfield et al. (2011); Stenger et al. (2023) . Simulating the dynamics of open systems, however, is less straightforward and demands a strategy to implement non-reversible operations.
@@ II.2 Open systems
Consider an open system of electrons coupled to metallic contacts. An example of such a system is a chain of electrons between source and drain contacts as illustrated in Fig. 1a. We represent each site by a qubit, where the end qubits are connected to the contacts as shown in Fig. 1b. Transport is modeled by evolution of the system state, in which electrons are injected or removed at the contacts. To simulate the dynamics we introduce the state |$\ket{\Phi(t)}$ at time $t$, starting from the initial state |$\ket{\Phi(0)}=\ket{\Phi_{0}}$.
@@
At each end site and time-point, we probabilistically select between three possible actions: (1) inject an electron, (2) remove an electron, (3) do nothing. Consider a quantum register with $L$ qubits and let |$\ket{0}_{q}$ and |$\ket{1}_{q}$ be the basis states for Hilbert space of qubit $q$. The actions are performed through mid-circuit measurements. An electron is injected by setting the qubit measured in |$\ket{0}_{q}$ to |$\ket{1}_{q}$ and removed by setting the qubit measured in |$\ket{1}_{q}$ to |$\ket{0}_{q}$. For each end qubit $q$ and conductor $\alpha$, we define
@@
where $P^{\text{in}}_{q\alpha}$ and $P^{\text{out}}_{q\alpha}$ are the probabilities of electron injection and removal, respectively, $f$ is the Fermi–Dirac distribution function, $\mu_{\alpha}$ is the electrochemical potential of contact $\alpha$, and $\eta_{q\alpha}=\Gamma_{q\alpha}t/N_{t}$, where $\Gamma_{q\alpha}$ is the tunneling rate Maslova et al. (2021); Liu et al. (2025) between qubit $q$ and conductor $\alpha$. Note that $N_{t}$ must be sufficiently large such that the probability of electron injection or removal is less than or equal to one, $N_{t}\geq(\Gamma_{q\alpha}t)^{-1}$ for all $\Gamma_{q\alpha}$. In the limit $N_{t}\rightarrow\infty$, this method simulates the dynamics of a system in contact with two conductors. See Appendix A for a comparison of our method with the Lindblad equation Brasil et al. (2013).
@@ III.1 Model System
We now demonstrate our method for the chain of interacting electrons, described by the Hamiltonian,
$\hat{H}=\gamma\sum_{<i,j>}\sum_{\sigma}(\hat{c}^{\dagger}_{i\sigma}\hat{c}_{j\sigma}+\hat{c}^{\dagger}_{j\sigma}\hat{c}_{i\sigma})+v\sum_{<i,j>}\sum_{\sigma}\hat{n}_{i\sigma}\hat{n}_{j\sigma},$ (4)
@@
$\begin{split}\hat{H}=\frac{\gamma}{2}\sum_{<q,p>}(\hat{X}_{q}\hat{X}_{p}+\hat{Y}_{q}\hat{Y}_{p})+\frac{v}{4}\sum_{<q,p>}(\hat{I}-\hat{Z}_{q})(\hat{I}-\hat{Z}_{p}).\\ \end{split}$ (5)
@@
We then simulate the electron chain both as a closed system, and as an open system coupled to contacts. In the closed-system case, we set $\eta_{q\alpha}=0$ for all $q$ and $\alpha\in\{\text{S},\text{D}\}$. In the open-system case, we couple the system to a source $f(\mu_{\text{S}})=1$ at the first site $\eta_{1\text{S}}\neq 0$ and a drain $f(\mu_{\text{D}})=0$ at last site $\eta_{L\text{D}}\neq 0$, where $i=L$ is the last site.
@@ III.3 Open-system dynamics
We now compare the dynamics of the open system simulated on a classical computer to those simulated on a quantum computer. Figure 3 shows the dynamics of a seven-site system as simulated on a classical (a-b) and quantum(c-d) computer. We consider $\gamma$=3.0 meV and 5.0 meV, and set $v$=10.0 meV, $\eta_{\text{qc}}=0.5$ meV, and $N_{t}/t=2\leavevmode\nobreak\ \text{meV}/\hbar$, in all cases. For all simulations shown in Fig.  3, the density is initially $n_{0}(0)=1.0$ and corresponds to the injection of an electron into the chain. Over time, the densitys at sites $i>1$ increase, as electron density flows towards the drain as indicated by the dotted green line in Fig. 3a. Black asterisks at $i=1$ correspond to the injection of an electron while white asterisks at $i=L$ correspond to the removal of electrons. We see that injecting an electron causes a sharp increase in density while the removal of an electron causes a sharp decrease in density, as expected. When electrons are not immediately removed, they can reflect off of the right end of the wire. Backpropagation of electron density can be seen as marked by the dotted yellow line. This results in a peak in density at the left end not corresponding to the injection of an electron. The flow of density from $i=1$ to $i=7$ is faster for $\gamma$=5.0 meV (Fig. 3b,d) than for $\gamma$=3.0 meV (Fig. 3a,c), on both the classical and quantum computers.
While the data taken from the quantum computer looks qualitatively very similar to that taken from the classical computer, there are errors arising from the quantum noise. In Table 2, we show selected data points from Fig. 3b,d. We see that the quantum computer tends to smooth the sharp peaks in electron density. This is likely due to decoherence from the quantum computer.
@@
Classical Quantum
$t\leavevmode\nobreak\ (\hbar/\text{meV})$ $n_{1}(t)$ $n_{4}(t)$ $n_{7}(t)$ $n_{1}(t)$ $n_{4}(t)$ $n_{7}(t)$
0.0 1.00 0.00 0.00 1.00 0.00 0.00
0.5 0.74 0.27 0.00 0.68 0.34 0.00
1.0 0.75 0.32 0.69 0.68 0.37 0.28
1.5 0.80 0.59 0.00 0.69 0.44 0.01
2.0 0.41 0.61 0.75 0.61 0.39 0.48
2.5 1.00 0.53 0.00 0.80 0.44 0.01
3.0 1.00 0.52 0.00 0.84 0.44 0.01
3.5 0.42 0.55 0.00 0.54 0.45 0.01
4.0 1.00 0.37 0.00 0.79 0.40 0.01
4.5 0.33 0.49 0.23 0.49 0.42 0.12
5.0 0.36 0.56 0.00 0.54 0.40 0.01
Table 2: Data from Fig. 3b (Classical) and Fig. 3d (Quantum).
@@ Appendix A Relationship to the Lindblad Equation
where $\hat{H}$ is the Hamiltonian of the system, $\hat{L}_{0}=\sqrt{\Gamma f(\mu)}\hat{c}^{\dagger}_{l}$ is the Lindblad operator for the incoming electrons, $\hat{L}_{1}=\sqrt{\Gamma[1-f(\mu)]}\hat{c}_{l}$ is the Lindblad operator for the outgoing electrons, $\Gamma$ is the tunneling rate between the contact and the system, $f(\mu)$ is the Fermi-Dirac distribution function evaluated at the chemical potential of the lead given that the lead has an electrochemical potential $\mu$, $\hat{c}^{\dagger}_{i}$ is the electron creation operator at point $i$, and $l$ is the point of contact. We use selective mid-circuit measurements to reproduce these dynamics on a quantum computer.
Figure 5: Simulations of open-system dynamics using Lindblad evolution (a) and our method (b), for $\gamma$=3.0 meV, and 5.0 meV. The source and drain are at $i=1$ and $i=7$, respectively with $P_{\text{in}}=P_{\text{out}}=0.5$. Both approaches use Trotterized time evolution with $N_{t}/t=2\leavevmode\nobreak\ \text{meV}/\hbar$
@@
We numerically compare our method to Lindblad evolution in Fig. 5 with select data being shown in Table 3. One striking difference is that, in the Lindblad evolution, the contacts are modeled as continuous injection or removal of fractional electron density. In our method the contacts probabilistically inject or remove a full electron. In the limit $\delta t\rightarrow 0$, these methods become indistinguishable, however, even for finite $\delta t$ the two methods share many qualitative features. Note that the density reaches the right edge at the same time in both methods, indication that the electron group velocity is the same for both methods. Furthermore, the average electron density in the center of the wire reaches the same value for both methods, at long times. It should be noted that Trotterization errors may accumulate differently for the two methods, especially because our method requires an extra approximation in order to convert the evolution operator into quantum gates.
Classical Lindblad
$t\leavevmode\nobreak\ (\hbar/\text{meV})$ $n_{1}(t)$ $n_{4}(t)$ $n_{7}(t)$ $n_{1}(t)$ $n_{4}(t)$ $n_{7}(t)$
4.0 1.00 0.51 0.00 0.87 0.50 0.09
4.1 1.00 0.50 0.00 0.88 0.50 0.09
4.2 1.00 0.54 0.12 0.88 0.50 0.09
4.3 0.90 0.55 0.00 0.88 0.51 0.09
4.4 0.67 0.52 0.00 0.88 0.51 0.09
4.5 0.49 0.49 0.12 0.89 0.52 0.09
4.6 1.00 0.51 0.38 0.89 0.52 0.09
4.7 1.00 0.56 0.58 0.89 0.52 0.09
4.8 1.00 0.52 0.00 0.89 0.53 0.09
4.9 0.89 0.46 0.00 0.89 0.53 0.09
5.0 0.64 0.48 0.00 0.89 0.53 0.09
5.1 0.45 0.44 0.00 0.89 0.53 0.09
5.2 0.35 0.57 0.13 0.90 0.54 0.09
5.3 1.00 0.58 0.41 0.90 0.54 0.09
5.4 0.88 0.55 0.59 0.90 0.54 0.08
5.5 1.00 0.46 0.64 0.90 0.54 0.08
5.6 1.00 0.41 0.66 0.90 0.55 0.08
5.7 0.89 0.53 0.00 0.90 0.55 0.08
5.8 0.63 0.64 0.00 0.90 0.55 0.08
5.9 0.41 0.62 0.12 0.90 0.55 0.08
6.0 0.33 0.59 0.38 0.90 0.55 0.08
avg. 0.78 0.53 0.20 0.89 0.53 0.09
Table 3: Data from Fig. 5a (Lindblad) and Fig. 5b (Classical).